Let f : R2-->R1be defined by
f(x1, x2) = (3x1+ 4x2)and T : R2--> R2be
defined by T(x1 , x2) = (x1- x2, x1+ x2).
Suppose g = f o T. What is g (2, 3) ?
1
Expert's answer
2020-06-01T19:19:09-0400

"f(x_1,x_2)=3x_1+4x_2\\
T(x_1,x_2)=(x_1-x_2,x_1+x_2)\\
g(x_1,x_2)=f(x_1,x_2)\circ T(x_1,x_2)=\\
=3(x_1-x_2)+4(x_1+x_2)=7x_1+x_2\\
g(2,3)=7\cdot2+3=17"